OpenStudy (anonymous):

Danielle is playing the newest app on her phone, Algebra City Defense. Earlier in the game, you and Danielle helped the Mathonauts defend the city against a giant, three-headed monster. After defeating the monster, the game advances to Level Two! You and Danielle must help the Mathonauts blast off to their Interstellar Headquarters. 1. Before blast-off, the Mathonauts must set the trajectory of their ship. Create a linear equation with a positive slope to be your original trajectory. @wolfe8

OpenStudy (anonymous):

I created a linear equation I just need help with a few of the questions.

OpenStudy (wolfe8):

What are the questions?

OpenStudy (anonymous):

Before departing the Interstellar Headquarters, the ship’s navigator begins to plan the next trip. She uses a map with their current trajectory already graphed on it. There is also a table that has coordinate points of a satellite that the Mathonauts must intercept. Explain to the navigator how she can use the graph and table to find where they will intercept the satellite. Assume the path of the satellite is linear. Use the coordinates to create an equation representing the path of the satellite. Explain the process and show how to find the point of intersection algebraically. The navigator has to program the ship’s computer so that it will be prepared to intercept the satellite. To program it, the satelite’s path is set equal to the ship’s. The programmed equation is –4x – 2 = –x + 3. Explain to the navigator how the programmed equation can be graphed to find the point of intersection. Use complete sentences.

OpenStudy (wolfe8):

You can separate them into 2 linear equations again as: y=-4x-2 and y=-x+3 Recall the general form y=mx+b where m is the slope and b is the y-intercept. Can you recognize m and b from the two equations?

OpenStudy (anonymous):

is this for number 5 or 6?

OpenStudy (wolfe8):

6. For 5, I would try each coordinate to see if they lie on the graph. Substitute x and y and see if the left and right hand side agree with each other.

OpenStudy (anonymous):

oh ok, and no not really? would m be -2?

OpenStudy (wolfe8):

m is the constant that goes with x see :)

OpenStudy (anonymous):

oh shoot I was thinking about b lol I see

OpenStudy (anonymous):

so you would just solve ?

OpenStudy (wolfe8):

Oh you don't need to solve anything, just read off what's in front of x haha ALTERNATIVE I guess an easier approach would be to pick 2 points for x and substitute them into the equation and plot the graph :)

OpenStudy (anonymous):

I don't think graphing is necessary, this is a virtual school assignment and they don't give graphs on assignments lol

OpenStudy (wolfe8):

Oh right. Well, you would then just say, I suppose, that you try every point in the table and see which one is on the straight trajectory line

OpenStudy (anonymous):

I just don't understand how to put it in sentences.

OpenStudy (wolfe8):

I put it in a sentence for you xD Can you make it sound nicer?

OpenStudy (anonymous):

thanks :)

OpenStudy (wolfe8):

No problem. Welcome again. And sorry I haven't been too enthusiastic.

OpenStudy (anonymous):

Youve been enthusiastic enough! I made you my fan

OpenStudy (anonymous):

I'm your fan I meant ahaa

OpenStudy (wolfe8):

Oh but I'm usually better than this. I was just saying bye to someone so I was sad. Well thank you :) All the best.

OpenStudy (anonymous):

I could not tell a bit. And you're welcome :)

OpenStudy (anonymous):

swag